OpenStudy (hazmatpuppy):

Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

TheSmartOne (thesmartone):

parallel lines have the same slope \(y - y_1 = m(x- x_1)\) where m is the slope what is the slope of the equation they gave you?

OpenStudy (hazmatpuppy):

2/3

TheSmartOne (thesmartone):

\(y - y_1 = \frac{2}{3}(x- x_1)\) now plug in the point (1, -2)

OpenStudy (hazmatpuppy):

-2 - y1 = 2/3(1 - x1)

OpenStudy (hazmatpuppy):

I also have to put it in in the form of _ _ - _ _ _ _

TheSmartOne (thesmartone):

replace y1 with -2 not y with -2 replace x1 with 1 not x with 1 :) you need it in y = mx + b form?

OpenStudy (hazmatpuppy):

No, standard form. so it would be y+2 = 2/3(x-1)? And then I'd put that in standard form?

OpenStudy (hazmatpuppy):

Also I cannot put fractions or decimals.

TheSmartOne (thesmartone):

yes, so let's multiply both sides by 3 to get rid of the fraction

OpenStudy (skullpatrol):

^ good idea :-D

OpenStudy (hazmatpuppy):

so y+6 = 2(x-3) And then distributive properties y+6 = 2x - 6 Correct?

OpenStudy (skullpatrol):

Nope, y + 6 should be 3y + 6, right?

TheSmartOne (thesmartone):

when you multiply 3 on one side, you have to do it on the other side too

OpenStudy (hazmatpuppy):

Right, I was thinking alike properties, so 3y + 6 = 6x - 6?

TheSmartOne (thesmartone):

I will be right back. I'm leaving you in good hands. :)

OpenStudy (hazmatpuppy):

Alrighty.

OpenStudy (skullpatrol):

Yes, now put it into standard form: ax + by = c where a, b, and c are integers.

OpenStudy (hazmatpuppy):

What should I do with the +6 and the -6?

OpenStudy (hazmatpuppy):

Can both be C?

OpenStudy (skullpatrol):

Yes. Combine them on one side.

OpenStudy (hazmatpuppy):

So 0?

OpenStudy (skullpatrol):

Nope. Subtract 6 from both sides.

OpenStudy (hazmatpuppy):

Not sure why I thought they'd cancel each other out. So I would get -12?

OpenStudy (skullpatrol):

Yes.

OpenStudy (hazmatpuppy):

So then I'd have to subtract 3y so I could get 6x - 3y = -12?

TheSmartOne (thesmartone):

3y + 6 = 6x - 6? We have to look at the big picture. :) If we subtract 6 on both sides 3y + 6 - 6 = 6x - 6 - 6 3y = 6x - 12 If you bring 3y on to the other side, it's not getting any closer to Standard Form. You could bring 6x to the other side and then multiply both sides by -1. Or you can bring 3y to the other side and then take -12 to the other side by adding 12 on both sides.

OpenStudy (skullpatrol):

3y + 6 = 6x - 6 3y + 6 - 6 = 6x - 6 - 6 3y + 0 = 6x -12 3y = 6x -12

OpenStudy (skullpatrol):

Standard form $$\Huge ax + by = c$$ where a, b and c are integers.

OpenStudy (hazmatpuppy):

Ah right, I'm doing this all n my head which is never good. But I don't have anything to write on besides notepad. So, subtract 6x to get -6x + 3y = -12 But since x is supposed to be a non-negatie it would be 6x - 3y = 12 Correct?

TheSmartOne (thesmartone):

That's why I suggested multiplying both sides by -1 Think of it like this a + b is the same as b + a (-6x) + (3y) is the same as (3y) - (6x) which would be 3y - 6x But we want it to be Ax + By So multiply both sides by -1 -6x + 3y = -12 Also, you can always type out your work on here and post it (:

OpenStudy (hazmatpuppy):

It doesn't fit, so should I divide by 2?

TheSmartOne (thesmartone):

Mhmm, 3 isn't divisible by 2

OpenStudy (hazmatpuppy):

Oh! Then 3.

OpenStudy (skullpatrol):

Yes. Divide both sides by 3.

OpenStudy (hazmatpuppy):

Alright, so now I have 2x - y = 6 The reason I have it as minus is because in the squares there's a forced subtraction symbol.

OpenStudy (hazmatpuppy):

Alright. I have 3 more, should I just open another discussion?

TheSmartOne (thesmartone):

You went from -6x + 3y = -12 to 6x - 3y = -12 We'll need to backtrack a bit :)

TheSmartOne (thesmartone):

Think of it like this a + b is the same as b + a (-6x) + (3y) is the same as (3y) - (6x) which would be 3y - 6x But we want it to be Ax + By So multiply both sides by -1 -1(-6x + 3y) = -1(-12 )

OpenStudy (hazmatpuppy):

Yes, I had 6x - 3y = 12 not -12

TheSmartOne (thesmartone):

Uh-oh We made a mistake in the beginning of the question. Earlier above, you changed it from (x-1) to (x-3) So starting from scratch: y+2 = 2/3(x-1) 3y + 6 = 2(x - 1) 3y + 6 = 2x - 2 3y + 8 = 2x And you can complete the last step :P

OpenStudy (hazmatpuppy):

What about when we multiplied everything by 3 to make 2/3 into 2?

OpenStudy (skullpatrol):

That only affects the 2/3

OpenStudy (skullpatrol):

Not the ( x - 1 )

TheSmartOne (thesmartone):

y+2 = 2/3(x-1) Multiply both sides by 3